# “Cold, Calculated, and Condescending”: How AI Identifies and Explains Ableism Compared to Disabled People

### v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXX
@@ 2. RQ2:
Our results highlight that SOTA models do not account for toxic speech targeted towards PwD, thereby leaving the internet unsafe for one-sixth of the global population. We contribute the first publicly available dataset of ableist speech (see link), and collaborate with PwD to assess how AI systems identify and explain disability bias, setting a pivotal foundation to bridging this gap. We discuss the challenges of benchmarking ableist speech, and provide concrete avenues for online ableism harm interventions. This effort calls for the active involvement of PwD in the AI development process, ensuring that intersectional disabled experiences are represented and prioritized.
@@ 3.1. Ableist Speech Dataset Curation
Given the absence of publicly available datasets for ableist speech, we assembled our own dataset of ableist and non-ableist sentences. In line with prior investigations of discriminatory experiences on social media (Brock, 2012; McGregor and Mourão, 2016; Farrell et al., 2019; Yang and Counts, 2018), we used Twitter and Reddit to search for explicit mentions or call-outs to ableist speech by people with disabilities. We used hashtags and keywords such as ”offensive,” ”ableist,” ”hate,” ”harassment,” ”problematic,” and ”discrimination,” and manually filtered for posts that targeted users with disabilities. To represent a diverse range of disabilities, as put forth by the Individuals with Disabilities Education Act (IDEA) (of Education, 2003), we also searched for these keywords in specific subreddit communities, such as r/autism, r/ADHD, r/blind, r/physicaldisabilities, and r/deaf, among others. We limited our date range to 2019-2024 to capture more recent instances of ableism. In addition to social media data, we also gathered primary accounts by PwD as reported in academic papers on ableist speech (Sannon et al., 2023; Heung et al., 2022, 2024; Lyu et al., 2024). We also gathered some non-ableist comments to assess for false positives, selecting toxic comments that were either unrelated to disability, (i.e., ”Reddit is so toxic sometimes!”), or comments that shared personal disability-related experiences, (i.e., “i can’t wear masks due to a cardiac issue”). In total, we collected 25 non-ableist comments in addition to 175 ableist comments, creating a comprehensive dataset of 200 comments.
Next, our team labeled each comment using three criteria: the specific disability it targeted (e.g., Blind or Low Vision, Learning Disability, Chronic Pain, or General if unspecified), the type of ableist speech using established taxonomies (Heung et al., 2022, 2024; Keller and Galgay, 2010), and whether the ableist bias was implicit or explicit. We coded comments with implicit bias when ableism appeared unintentionally in the form of microaggressions, invasive questions, or patronizing sentiments (e.g., “Wow, you’re so brave for living with a disability.”). In contrast, comments with explicit bias were intentionally hateful and manifested as slurs, threats, or eugenics-related speech (e.g., “You are the r-word.”)  (Venkit et al., 2022; Friedman and Gordon, 2023). An moderate Inter-Rater Reliability score (kappa = 0.59) depicts consensus of implicit and explicit bias rating within team members. Overall, our dataset represented 9 different disabilities, 10 types of ableist speech, and a mix of implicit bias (n=113, 65%) and explicit bias (n=62, 35%). This dataset is made publicly available here.
@@ 3.2. Obtaining Harm Assessments
From Toxicity Classifiers and LLMs. We used a Python script to collect toxicity ratings for each comment from three state-of-the-art (SOTA) toxicity classifiers: Google Jigsaw’s Perspective API (API, 2024), Azure’s AI Content Safety API (Safety, 2024), and OpenAI’s Moderation API (Guide, 2024). Each toxicity classifier returned various toxicity values (e.g., Perspective gave attributes for Profanity, Sexually Explicit Content, and Threats); we chose the maximum value among all returned values for a comment to account for any and all harm. Next, we prompted four SOTA LLMs—GPT-4 Turbo, Gemini Flash, Claude 3.5 Sonnet, and Llama 3.1—to evaluate each comment for toxicity and ableism on a scale of 1-10, and provide a rationale for their ableism scores. We used zero-shot prompting and asked models to evaluate harm in a grounded role-playing scenario, as exemplified by prior work on zero-shot reasoning (Kong et al., 2023) (see Appendix A.2 for full prompt).
A common concern with LLM output is its sensitivity to prompt engineering (White et al., 2023). To account for this, we ran a total of 20 trials, for all 200 comments and all 4 models. In these trials, we randomized the order of the comments, and attempted multiple variations of the prompt, such as asking it to ”think step by step” (Wei et al., 2022) for its ableism ratings and explanations. We noticed no significant difference in output, likely due to the nature of the task, which asks to return numerical ratings and short text responses rather than long-form text. We then used averaged toxicity and ableism scores, and randomly selected an explanation for each comment across all trials.
From PwD and Non-PwD. To assess alignment between AI models and PwD, we gathered ratings and explanations of toxicity and ableism from 130 PwD. We also collected ratings from 60 Non-PwD to examine whether LLMs carry normative tendencies and align closer with non-disabled ratings. All participants were recruited and screened using Prolific; eligible participants were directed to our Qualtrics survey. Each participant evaluated up to 10 comments from the dataset, rating toxicity and ableism on a 10-point Likert scale (Preston and Colman, 2000) (see Appendix A.1 for Qualtrics Survey). A subset of participants were asked to justify their scores, explaining why comments were (or were not) ableist. Overall, we collected 2,600 harm ratings—1,300 toxicity ratings and 1,300 ableism ratings—and over 800 explanations of ableism.
@@ 3.3. Assessing Alignment
Qualitative Analysis.  We conducted a study with 40 PwD to understand their perceptions of ableism explanations by LLMs. We selected a random subset of comments (n=25) and explanations (n=100) to review, to allow for in-depth qualitative analysis. Evaluations were gathered via two surveys. The first survey displayed the social media comment followed by one LLM explanation, where PwD assessed the quality of the explanation and suggested improvements. The second survey displayed the social media comment and two randomly chosen LLM explanations, where PwD selected their preferred explanation and explained their reasoning. This comparison focused on GPT-4 and Gemini responses, as their explanations differed significantly in length and quality, allowing us to examine how these factors influenced participant preferences.
@@ 3.4. Participant Demographics
Participant ages ranged from 18 to 74 (mean = 38.2, SD = 12.0), and 88 identified as men (46%), 97 as women (51%), and 5 preferred not to disclose (3%). Majority self-described as White (n = 134, 70%), while 16 (8%) identified as Black, 14 (7%) identified as Asian, and 14 (7%) identified as mixed race. All participants lived in the United States and reported native English fluency. Among the 130 PwD, 75 reported psychological disorders (i.e., Attention Deficit, Emotional Disturbance), 49 had orthopedic/mobility impairments (i.e., Cerebral Palsy, neuromotor disabilities), 42 were blind and/or low vision, 35 were on the Autism Spectrum, 27 had learning disabilities (i.e., Dyslexia), and 20 were being d/Deaf or Hard of Hearing. Many participants (n=73) reported more than one disability.
@@ 4.2. Human-AI Alignment in Ableism Identification (RQ1)
Normative Tendencies. We next examined if there are normative tendencies in how LLMs rate ableism. We found that Non-PwD rated ableism significantly lower than PwD $(Z=-2.86,p<0.01,r=0.29)$. Looking at effect sizes in pairwise Wilcoxon Signed-Rank tests, we found that ratings from Claude, Llama, and GPT-4 were more aligned with PwD, whereas from Gemini were more aligned with Non-PwD. For example, Gemini and PwD $(Z=-6.90,p<0.01,r=0.49)$ had a larger effect size, compared to Gemini and Non-PwD $(Z=-2.73,p<0.01,r=0.19)$, indicating a closer proximity to Non-PwD. These findings highlight that while some LLMs like Claude and Llama are more aligned with PwD in how they rate ableism, others like Gemini need more calibration to better reflect perspectives of PwD (see Appendix 5)
@@ 4.3.4. LLM explanations were condescending and judgmental, instead of educational
For instance, for the comment, “Autism is so over-diagnosed these days,” LLMs offered general explanations, such as:
GPT-4: ‘‘[this comment] perpetuates stigma and doubt against autistic people.’’
Claude: ‘‘[this comment] undermines the real experiences of autistic individuals.’’
In contrast, PwD provided richer context, framing the statement as part of a historical pattern of misconceptions:
@@ 4.3.5. PwD drew connections with other marginalized identities
“Analogously, most people are cisgender and heterosexual, but that does not mean we should pathologise those who are queer. Just as most struggles faced by queer people come from societal barriers favoring cisgender heterosexual people, most struggles faced by autistic people come from societal barriers favoring neurotypical people.”
@@ 5. Discussion
Online platforms are inundated with toxic speech towards people with disabilities—a historically marginalized group that makes up nearly one-sixth of the world’s population and endures disproportionate amounts of hate and bias (Organization, 2005; Ringland, 2019; Kaur et al., 2024; Heung et al., 2024; Sannon et al., 2023). AI models, that are often called to moderate this online harm, are riddled with biases against people with disabilities, perpetuating the very harm they aim to prevent (Venkit et al., 2022; Hassan et al., 2021; Hutchinson et al., 2020). Our work builds on this growing body of research and reveals a critical limitation: tbfAI models con, designed to curb harmful content,sistently underestimated toxicity compared to PwD,people with disabilitiesiland offered inconsistent assessments of ableismṪheir anations wereof ableism flawfundamentally ed—they lacked nuance, maderelied onorrect assumptions, and appeoften ared judgmental. Building on discourse around disability and AI fairness (Whittaker et al., 2019; Trewin, 2018; Bennett and Keyes, 2020), we discuss pervasive challenges in identifying ableist language, and propose actionable pathways for mitigating the harms of online ableism.
@@
Challenges of Identifying Ableist Speech.  Our dataset covers targeted language towards nine disabilities, including implicit and explicit language. While this dataset exposes critical harms in how AI models identify and explain ableism, it merely scratches the surface of disability representation. Developing a benchmarking dataset for ableist speech carries significant risks of representational harm—risks that are especially pronounced for vulnerable populations (Blodgett and O’Connor, 2017; Blodgett et al., 2020; Selvam et al., 2023; Rauh et al., 2022). Disabilities present unique representational challenges due to their complex and dynamic nature. Many are invisible, progressive, or exist along a spectrum of ability (e.g., Autism Spectrum Disorder, Vision Impairment). These nuances make it challenging to represent disabilities using static, binary markers (e.g., “blind,” “not blind”). Also, disability identities often overlap; 56% of participants in our study identified as having more than one disability. This complexity raises a critical question: how can we create a dataset for ableist speech that meaningfully represents these diverse, yet overlapping disability identities?
@@
Disability scholars recognize the complexity of identifying ableist speech, emphasizing that harm may stem from unconscious and implicit biases embedded within language (Milton, 2020). For example, phrases once considered acceptable are deemed ableist (e.g., “it fell on deaf ears”), while terminology that was previously ableist may be reclaimed, such as the shift from person-first to identity-first language (Dunn and Andrews, 2015; Sharif et al., 2022). Adding to this challenge, ableism is inherently subjective—what one perceives as highly ableist may be viewed as less harmful by another—and our study found some variance in how disabled participants rated ableism. Researchers highlight the importance of embracing this diversity to capture nuanced lived experiences within AI models (Wang et al., 2023; Plank, 2022; Prabhakaran et al., 2024; Chen et al., 2023; Parrish et al., 2024), presenting a critical challenge towards benchmarking ableist speech: how can we create representative datasets that capture the majority perspective on ableism, while also accounting for and honoring the unique experiences of minoritized disabled individuals?
@@
Explanations to Mitigate Ableism. The systemic flaws in AI models place an undue burden on disabled users, forcing them to report ableist speech and take on the responsibility of maintaining safe online spaces (Heung et al., 2022). Disability scholars have long advocated for a societal reckoning with ableism, emphasizing the shared responsibility of non-disabled people to confront and dismantle disability bias (Nikki, 2022). This important effort requires non-disabled people to reshape their perceptions of disability, deepen their understanding of ableism, and adopt inclusive language and behaviors. This raises a pressing question: Can AI systems encourage conscientious communication among non-disabled users? If so, how?

### v3

@@ Abstract.
large language models, disability, ableism, natural language, toxicity classifiers, hate speech
journalyear: 2025copyright: acmlicensedconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732128isbn: 979-8-4007-1482-5/2025/06ccs: Social and professional topics People with disabilitiesccs: Computing methodologies Artificial intelligence
Figure 1. A representation of AI models being unable to identify and explain ableism, as it relates to lived disability experiences.
[…]
A diagram comparing how different evaluators: Large Language Models (LLMs), Toxicity Classifiers (TCs), People with disabilities (PwD), and People without disabilities—rate and explain harm from a dataset comment: ’It’s survival of the fittest.’ The flow begins with ’Dataset Comments’ on the left, showing colored dots representing various comments, leading to four evaluator categories. On the right, there are different harm measurements. PwD assess the comment as ’Not ableist, but toxic,’ with a 20% harm rating, explaining that ’Fetishizing people is always wrong.’ TCs provide a 38% toxicity score. LLMs rate the comment as 60% harmful, stating ’This is significantly ableist.’
@@ 2. RQ2:
Our results highlight that SOTA models do not account for toxic speech targeted towards PwD, thereby leaving the internet unsafe for one-sixth of the global population. Our focus on ableism explanations emphasizes how biased AI assessments could misrepresent disability narratives, making tone, language, and verbiage critical in addressing disability bias. We contribute the first publicly available dataset of ableist speech (see link), and collaborate with PwD to assess how AI systems identify and explain disability bias, setting a pivotal foundation to bridging this gap. We discuss the challenges of benchmarking ableist speech, and provide concrete avenues for online ableism harm interventions. This effort calls for the active involvement of PwD in the AI development process, ensuring that disabled experiences are represented and prioritized.
@@ 3.1. Ableist Speech Dataset Curation
Gathering Ableist and Non-Ableist Data. Given the absence of publicly available datasets for ableist speech, we assembled our own dataset of ableist and non-ableist sentences. In line with prior investigations of discriminatory experiences on social media (Brock, 2012; McGregor and Mourão, 2016; Farrell et al., 2019; Yang and Counts, 2018), we used Twitter and Reddit to search for explicit mentions or call-outs to ableist speech by people with disabilities. We used hashtags and keywords such as ”offensive,” ”ableist,” ”hate,” ”harassment,” ”problematic,” and ”discrimination,” and manually filtered for posts that targeted users with disabilities. To represent a diverse range of disabilities, as put forth by the Individuals with Disabilities Education Act (IDEA) (of Education, 2003), we also searched for these keywords in specific subreddit communities, such as r/autism, r/ADHD, r/blind, r/physicaldisabilities, and r/deaf, among others. We limited our date range to 2019-2024 to capture more recent instances of ableism. In addition to social media data, we also gathered primary accounts by PwD as reported in academic papers on ableist speech (Sannon et al., 2023; Heung et al., 2022, 2024; Lyu et al., 2024).
We also gathered some non-ableist comments to assess for false positives, in line with benchmarking practices that require representations from both binaries of an attribute to detect disparities (e.g., ableist and non-ableist), for example, #DisabledOnIndiaTwitter (Mondal et al., 2022). We selected social media comments that were either unrelated to disability, (i.e., ”Reddit is so toxic sometimes!”), or comments that shared personal disability-related experiences, (i.e., “i can’t wear masks due to a cardiac issue”). In total, we collected 25 non-ableist comments in addition to 175 ableist comments, creating a comprehensive dataset of 200 comments.
Labeling Data. To ensure broad coverage in our dataset, we labelled all comments with (1) the specific disability it targeted (e.g., Blind or Low Vision, Autism, Chronic pain, or General if unspecified) (of Education, 2003), (2) the type of ableist speech (e.g., slurs & derogatory language, patronization, etc.) (Heung et al., 2022, 2024; Keller and Galgay, 2010) and (3) the implicit or explicit nature of the ableist bias. We coded comments with implicit bias when ableism appeared unintentionally in the form of microaggressions, invasive questions, or patronizing sentiments (Keller and Galgay, 2010; Heung et al., 2022; Venkit et al., 2022) (e.g., “Wow, you’re so brave for living with a disability.”). In contrast, comments with explicit bias were intentionally hateful and manifested as slurs, threats, or eugenics-related speech (e.g., “You are the r-word.”)  (Venkit et al., 2022; Friedman and Gordon, 2023). A moderate Inter-Rater Reliability score (kappa = 0.59) depicts consensus of implicit and explicit bias rating within team members. Overall, our dataset represented 9 different disabilities, 10 types of ableist speech, and a mix of implicit bias (n=113, 65%) and explicit bias (n=62, 35%). This dataset is made publicly available here: link.
@@ 3.2. Obtaining Harm Assessments
From Toxicity Classifiers and LLMs. We used a Python script to collect toxicity ratings for each comment from three state-of-the-art (SOTA) toxicity classifiers: Google Jigsaw’s Perspective API (API, 2024), Azure’s AI Content Safety API (Safety, 2024), and OpenAI’s Moderation API (Guide, 2024). Each toxicity classifier returned various toxicity values (e.g., Perspective gave attributes for Profanity, Sexually Explicit Content, and Threats); we chose the maximum value among all returned values for a comment to account for any and all harm. Next, we prompted four SOTA LLMs—GPT-4 Turbo, Gemini Flash, Claude 3.5 Sonnet, and Llama 3.1—to evaluate each comment for toxicity and ableism on a scale of 1-10, and provide a rationale for their ableism scores. We used zero-shot prompting and asked models to evaluate harm in a grounded role-playing scenario, as exemplified by prior work on zero-shot reasoning (Kong et al., 2023) (see Appendix A.2 for full prompt). To further assess baseline performance of these models, we did not include definitions of toxicity and ableism in our prompt.
A common concern with LLM output is its sensitivity to prompt engineering (White et al., 2023). To account for this, we ran a total of 20 trials, for all 200 comments and all 4 models. In these trials, we randomized the order of the comments, and attempted multiple variations of the prompt, such as asking it to ”think step by step” (Wei et al., 2022) for its ableism ratings and explanations. We noticed no significant difference in output, likely due to the nature of the task, which asks to return numerical ratings and short text responses rather than long-form text. We then averaged toxicity and ableism scores, and randomly selected an explanation for each comment across all trials.
From PwD and Non-PwD. To assess alignment between AI models and PwD, we gathered ratings and explanations of toxicity and ableism from 130 PwD. We also collected ratings from 60 Non-PwD to examine whether LLMs carry normative tendencies and align closer with non-disabled ratings. All participants were recruited and screened using Prolific; eligible participants were directed to our Qualtrics survey.
Each participant evaluated up to 10 comments from the dataset, rating toxicity and ableism on a 10-point Likert scale (Preston and Colman, 2000) (see Appendix A.1 for Qualtrics Survey). We borrow the term “toxic” from Kumar et al’s (Kumar et al., 2021) study which surveyed 300 diverse participants with adjacent terms like “hateful,” or “offensive,” finding that “toxic” encompassed the largest umbrella of hate. Similar to Perspective’s instructions for annotators, we provided participants with definitions of toxicity and ableism. Our definition of toxicity was akin to Perspective’s, “A rude, disrespectful, or unreasonable comment that is likely to make people leave a discussion.” (API, 2024). Our definition of ableism was derived from Keller and Galgay’s (Keller and Galgay, 2010) foundational work on microaggressive ableism, and Nelson’s Ableism chapter in the Handbook of Prejudice, Stereotyping, and Discrimination (Nelson, 2024).
A subset of participants were asked to justify their scores, explaining why comments were (or were not) ableist. Overall, we collected 2,600 harm ratings—1,300 toxicity ratings and 1,300 ableism ratings—and over 800 explanations of ableism.
@@ 3.3. Assessing Alignment
Qualitative Analysis.  We conducted an independent study with 40 PwD to understand their perceptions of ableism explanations by LLMs. We selected a random subset of comments (n=25) and explanations (n=100) to review, to allow for in-depth qualitative analysis. Evaluations were gathered via two surveys. The first survey displayed the social media comment followed by one LLM explanation, where PwD assessed the quality of the explanation and suggested improvements. The second survey displayed the social media comment and two randomly chosen LLM explanations, where PwD selected their preferred explanation and explained their reasoning. This comparison focused on GPT-4 and Gemini responses, as their explanations differed significantly in length and quality, allowing us to examine how these factors influenced participant preferences.
@@ 3.4. Participant Demographics
We present the demographics of all 170 PwD that participated in our study, 130 PwD which annotated toxicity and ableism, and an additional 40 PwD that evaluated AI annotations.
Participant ages ranged from 18 to 74 (mean = 38.2, SD = 12.0), and 78 identified as men (46%), 87 as women (51%), and 5 preferred not to disclose (3%). Majority self-described as White (n = 134, 70%), while 16 (8%) identified as Black, 14 (7%) identified as Asian, and 6 (4%) identified as mixed race. All participants lived in the United States and reported native English fluency. Among the 170 PwD, 98 (58%) reported psychological disorders (i.e., Attention Deficit, Emotional Disturbance), 64 (38%) had orthopedic/mobility impairments (i.e., Cerebral Palsy, neuromotor disabilities), 55 (32%) were blind and/or low vision, 46 (27%) were on the Autism Spectrum, 35 (21%) had learning disabilities (i.e., Dyslexia), and 26 (15%) were being d/Deaf or Hard of Hearing. Many participants (n=95) reported more than one disability.
@@ 4.2. Human-AI Alignment in Ableism Identification (RQ1)
Normative Tendencies. We next examined if there are normative tendencies in how LLMs rate ableism. We found that Non-PwD rated ableism significantly lower than PwD $(Z=-2.86,p<0.01,r=0.29)$. Looking at effect sizes in pairwise Wilcoxon Signed-Rank tests, we found that ratings from Claude, Llama, and GPT-4 were more aligned with PwD, whereas from Gemini were more aligned with Non-PwD. For example, Gemini and PwD $(Z=-6.90,p<0.01,r=0.49)$ had a larger effect size, compared to Gemini and Non-PwD $(Z=-2.73,p<0.01,r=0.19)$, indicating a closer proximity to Non-PwD. These findings highlight that while some LLMs like Claude and Llama are more aligned with PwD in how they rate ableism, others like Gemini need more calibration to better reflect perspectives of PwD (see Appendix A.4)
@@ 4.3.4. LLM explanations were condescending and judgmental, instead of educational
For instance, for the comment, “Autism is so over-diagnosed these days,” LLMs offered general explanations, suggesting this comment, ‘‘perpetuates stigma and doubt against autistic people,’’ (GPT-4), or it ‘‘undermines the real experiences of autistic individuals’’ (Claude). In contrast, PwD provided richer context, framing the statement as part of a historical pattern of misconceptions:
@@ 4.3.5. PwD drew connections with other marginalized identities
“Most people are cisgender and heterosexual, but that does not mean we should pathologise those who are queer. Just as most struggles faced by queer people come from societal barriers favoring cisgender heterosexual people, most struggles faced by autistic people come from societal barriers favoring neurotypical people.”
@@ 5. Discussion
Online platforms are inundated with toxic speech towards people with disabilities—a historically marginalized group that makes up nearly one-sixth of the world’s population and endures disproportionate amounts of hate and bias (Organization, 2005; Ringland, 2019; Kaur et al., 2024; Heung et al., 2024; Sannon et al., 2023). AI models, that are often called to moderate this online harm, are riddled with biases against people with disabilities, perpetuating the very harm they aim to prevent (Venkit et al., 2022; Hassan et al., 2021; Hutchinson et al., 2020). Our work builds on this growing body of research and reveals a critical limitation: AI models, designed to curb harmful content, consistently underestimated toxicity compared to people with disabilities, and offered inconsistent assessments of ableism. Their explanations of ableism were fundamentally flawed—they lacked nuance, relied on incorrect assumptions, and often appeared judgmental. Building on discourse around disability and AI fairness (Whittaker et al., 2019; Trewin, 2018; Bennett and Keyes, 2020), we discuss pervasive challenges in identifying ableist language, and propose actionable pathways for mitigating the harms of online ableism.
Challenges of Identifying Ableist Speech.  Our study aligns with previous research on bias in toxicity detection, reaffirming that AI tools struggle to interpret implicit bias and context (Pavlopoulos et al., 2020; Han and Tsvetkov, 2020). Differently from these studies however, we center the voices of disabled people, exposing how automated content moderation can reinforce ableist harm, and reduce disability visibility. Despite being the largest minority group, and receiving copious amounts of hate online, toxicity classifiers abhorrently failed at identifying ableism. Even prior research on toxicity classification takes into account race, gender, sexuality (Kumar et al., 2021; Muralikumar et al., 2023), but not disability. This demographic continues to be excluded from discourse and development of automated content moderation.
Our dataset covers targeted language towards nine disabilities, including implicit and explicit language. While this dataset exposes critical harms in how AI models identify and explain ableism, it merely scratches the surface of disability representation. Developing a benchmarking dataset for ableist speech carries significant risks of representational harm—risks that are especially pronounced for vulnerable populations (Blodgett and O’Connor, 2017; Blodgett et al., 2020; Selvam et al., 2023; Rauh et al., 2022). Ableism is a byproduct of disabling environments (Oliver, 2013), and can manifest online variably through multiple means, such as platform-mediated harm (e.g., a person with dwarfism being misclassified as a minor on TikTok (Heung et al., 2022)), or language embedded with unconscious biases (e.g., “You don’t look autistic.”). This is a call to expand our definition of online toxicity, to include these direct and indirect forms of ableism.
@@
Moreover, disabilities present unique representational challenges due to their complex and dynamic nature. Many are invisible, progressive, or exist along a spectrum of ability (e.g., Autism Spectrum Disorder, Vision Impairment). These nuances make it challenging to represent disabilities using static, binary markers (e.g., “blind,” “not blind”). Also, disability identities often overlap; 56% of participants in our study identified as having more than one disability. This complexity raises a critical question: how can we create a dataset for ableist speech that meaningfully represents these diverse, yet overlapping disability identities?
@@
Disability scholars recognize the complexity of identifying ableist speech, emphasizing that harm may stem from unconscious and implicit biases embedded within language (Milton, 2020). For example, phrases once considered acceptable are deemed ableist (e.g., “it fell on deaf ears”), while terminology that was previously ableist may be reclaimed, such as the shift from person-first to identity-first language (Dunn and Andrews, 2015; Sharif et al., 2022). Adding to this challenge, ableism is inherently subjective—what one perceives as highly ableist may be viewed as less harmful by another—and our study found variance in how disabled participants rated ableism. Researchers highlight the importance of embracing this diversity to capture nuanced lived experiences within AI models (Wang et al., 2023; Plank, 2022; Prabhakaran et al., 2024; Chen et al., 2023; Parrish et al., 2024), presenting a critical challenge towards benchmarking ableist speech: how can we create representative datasets that capture the majority perspective on ableism, while also accounting for and honoring the unique experiences of minoritized disabled individuals?
@@
Explanations to Mitigate Ableism. The systemic flaws in AI models place an undue burden on disabled users, forcing them to report ableist speech and take on the responsibility of maintaining safe online spaces (Heung et al., 2022). Disability scholars have long advocated for a societal reckoning with ableism, emphasizing the shared responsibility of non-disabled people to confront and dismantle disability bias (Nikki, 2022; Branco et al., 2019; Cencirulo et al., 2021). This important effort requires non-disabled people to reshape their perceptions of disability, deepen their understanding of ableism, and adopt inclusive language and behaviors. This raises a pressing question: Can AI systems encourage conscientious communication among non-disabled users? If so, how?